OpenStudy (anonymous):

How many grams of hydrogen gas must be "burned" to form 60.0 grams of water?

OpenStudy (abb0t):

Are you familiar with combustion reaction?

OpenStudy (anonymous):

\[H _{2}+\frac{ 1 }{ 2 }O _{2}\rightarrow H _{2}O\] so according to the reaction 2*(1.01) grams of hydrogen makes 2*(1.01)+15.999 = 18.019 grams of water so to make 60 grams of water amount of hydrogen required will be equal to (2.02*60)/18.019 = 6.73 grams